OpenStudy (anonymous):

USe the function to evaluate the indicated expressions and simplify f(x)=x+4 f(x^2),(f(x))^2

OpenStudy (cgreenwade2000):

One says to replace x with x^2 in the function. The other one says to take the entire function and square it.

OpenStudy (anonymous):

f(x)=x+4 f(x^2)=x^2 +4 (f(x))^2=(x+4)^2 =x^2 +8x+16

OpenStudy (cgreenwade2000):

We're not really here to do other people's homework mark.

OpenStudy (cgreenwade2000):

Yes, shake your head at me because you learned how to do the problem. Oh wait!